OpenStudy (anonymous):

Determine the limit of the trigonometric function (if it exists). lim t->0 (sin3t)/(2t)

OpenStudy (anonymous):

Since this form is indeterminate you will have to use l'hospitals rule? Do you know how to use it?

OpenStudy (anonymous):

never heard of it.

OpenStudy (anonymous):

lol you need it!

OpenStudy (anonymous):

No, you don't have to use l'Hospital's. I don't recommend it in limits without knowing derivatives.

OpenStudy (anonymous):

We define \(u=\frac{1}{3}t\) We know that as \(t\to0\), \(u\to 0\), so we can say: \[ \lim_{t\to0}\left(\frac{\sin(3t)}{2t}\right)=\lim_{u\to0}\left(\frac{\sin(u)}{\frac{2}{3}u}\right) \]We know, \(\lim_{x\to0}\frac{\sin(x)}{x}=1\) So: \[ \lim_{u\to0}\left(\frac{\sin(u)}{\frac{2}{3}u}\right)=1\lim_{u\to0}\left(\frac{1}{\frac{2}{3}}\right)=\frac{3}{2} \]Et voilá. Another nice way to do it, is using asymptotic notation: \[ \sin(x)\sim x, x\to0\]So, we substitute \(ax\) for \(\sin(ax)\) in the limit.

OpenStudy (anonymous):

And, my bad on the \(\frac{2}{3}u\), it's supposed to be \(2u\).